# On priors and scale cuts in EFT-based full-shape analyses

### v1

@@ Abstract
Parameter estimation from galaxy survey data from the full-shape method depends on scale cuts and priors on EFT parameters. The effects of priors, including the so-called “prior volume” phenomenon have been originally studied in Ivanov et al. (2019) and subsequent works. In this note, we repeat and extend these tests and also apply them to the priors used by D’Amico et al. (2019). We point out that in addition to the “prior volume” effect there is a much more dangerous effect that is largely overlooked: a systematic bias on cosmological parameters due to overoptimistic scale cuts. Unlike the “prior volume” effect, this is a genuine systematic bias due to two-loop corrections that does not vanish with better priors or with larger data volumes. Our study is based on the high fidelity BOSS-like PT Challenge simulation data which offer many advantages over analyses based on synthetic data generated with fitting pipelines. We show that the analysis choices of D’Amico et al. (2019), especially their scale cuts, significantly bias parameter recovery, overestimating $\sigma_{8}$ by over $5\%$ (equivalent to $1\sigma$). The bias on measured EFT parameters is even more significant. In contrast, the analysis choices associated with the CLASS-PT code lead to a negligible ($\lesssim 1\%$) bias in the recovery of cosmological parameters based on their best-fit values.
@@ 1 Introduction
The presence of the theory systematic error in EFT-FS analyses shows why parameter projection effects should not be considered in isolation. For instance, one may choose overoptimistic scale cuts or priors in their analysis, which would result in a very small prior volume effect, but both 1D posteriors and the best-fit values will be biased away from the true values. As we show here, this is the case for the analysis choices associated with the PyBird code D’Amico et al. (2021).
@@
The use of high-fidelity BOSS-like PT Challenge simulation data offers many advantages over analyses of based on synthetic data generated within the fitting pipeline, done e.g. in Simon et al. (2023). First, it enables the estimation of the theory systematic bias arising from the imperfect modeling of the target statistics. Second, it allows one to estimate the quality of the recovery of true cosmological parameters. Third, the large-volume data allow one to easily estimate the true best-fit model by rescaling the covariance. This method is thus not sensitive to noise in the covariance matrices Hartlap et al. (2007); Sellentin and Heavens (2016); Howlett and Percival (2017); Wadekar et al. (2020); Philcox et al. (2021) and does not suffer from numerical uncertainties associated with the likelihood minimization.
@@ 2 Summary and Main Conclusions
We compare analysis choices associated with the PyBird and CLASS-PT EFT-FS likelihoods. Following Nishimichi et al. (2020) we dub those two analysis configurations as “West coast” (WC) and “East coast” (EC) models. We also consider two extensions of these models, EC1/2 designed to highlight certain analysis aspects. The EC model that matches the actual BOSS analysis Philcox and Ivanov (2022) is referred to as EC3. Note that the WC analysis here is the one was used in the actual BOSS analysis D’Amico et al. (2019). [^6] It is different from the WC analysis implemented in the PT Challenge Nishimichi et al. (2020), which used a more conservative $k_{\mathrm{max}}$ value.
@@
The recovered best-fit value of $\sigma_{8}$ is shifted up by $\approx 5\%$ as a result of the two-loop corrections. This bias remains in the combinations with other datasets like Planck CMB. This raises serious concerns on the robustness of results obtained under analyses choices of D’Amico et al. (2019).
@@
Finally, our single-volume results can be treated as an approximation for upcoming surveys such as DESI and Euclid, for which the volumes of individual data chunks will be comparable to the entire volume of BOSS DR12, c.f. Chudaykin and Ivanov (2019); Aghamousa et al. (2016). As we show here, in this situation the prior volume effects should not significantly complicate the interpretation of the results from these surveys.
@@ 3 Comparison of different analyses choices and priors
We now proceed to the description of the counterterms and stochasticity parameters. The exact relations for those coefficients mildly depend on redshift. In what follows we will focus on the $z_{3}$ redshift bin ($z_{\rm eff}=0.61$). For the $z_{1}$ redshift slice ($z_{\rm eff}=0.38$) we found the similar results which we will report in the end of the section. In our calculations we adopt the fiducial $\Lambda$CDM cosmology used in the official BOSS analysis Alam et al. (2017b): $\Omega_{m}=0.31$, $h=0.676$, $\omega_{b}=0.022$, $\sigma_{8}=0.8$ and $n_{s}=0.97$. We also adopt $b_{1}=2.2$ from the same BOSS analysis.
@@
where D’Amico et al. (2019) assumed $k_{M}=0.7\,h\text{Mpc}^{-1}$. Matching them, we find the following relation between $c_{0}$ and $c_{ct}$ parameters,
@@
where $k_{\rm NL}=0.45\,h\text{Mpc}^{-1}$, $\bar{n}_{\rm WC}$ and $\bar{n}_{\rm EC}$ refer the mean galaxy densities in the WC and EC models. Notice that WC assume $\bar{n}_{\rm WC}^{-1}=2500\,(h^{-1}\text{Mpc})^{3}$ and $2200\,(h^{-1}\text{Mpc})^{3}$ for the LOWZ and CMASS redshift bins, whereas in the EC pipeline $\bar{n}_{\rm EC}^{-1}=3500\,(h^{-1}\text{Mpc})^{3}$ and $5000\,(h^{-1}\text{Mpc})^{3}$ for $z_{1}$ and $z_{3}$ slices. Note that the actual measurements from data give $\bar{n}^{-1}=4800$ and $3500$ $(h^{-1}\text{Mpc})^{3}$ for CMASS and LOWZ, respectively Chen et al. (2024). The values of $\bar{n}$ for the $z_{1}/z_{3}$ data split in the EC pipeline correspond to the mean geometric number densities, i.e. the total number of galaxies divided by the effective geometric volume. These numbers are in good agreement with $\bar{n}(z_{\rm eff})$ measured directly from data using the appropriate weights and selection functions.
@@
Adopting $c_{\epsilon,\,\text{quad}}\sim\mathcal{N}(0,2^{2})$ from D’Amico et al. (2019), we observe that the WC prior on $a_{2}$ is tighter than that in the EC model which is $a_{2}\sim\mathcal{N}(0,1^{2})$. The inconsistency of this choice with the physics of the HOD-based galaxies was discussed in Ivanov et al. (2024b). As we will see later, this tight prior reduces the parameter shifts caused by marginalization effects in a multi-chunk full-shape EFT analysis. Importantly, the WC setup also excludes a scale-dependent stochastic counterterm in the monopole, $c_{\epsilon,\,\text{mono}}=0$. Although the real-space scale-dependent noise term can be set to zero based on the simulation results Schmittfull et al. (2019, 2020); Ivanov et al. (2024a, b), the contribution from the $k^{2}\mu^{2}$ counterterm is actually non-negligible, which raises concerns about the choice $c_{\epsilon,\,\text{mono}}=0$.
@@
$P^{\rm stoch}(k,\mu)=\frac{1}{\bar{n}}\left(P_{\rm shot}+a_{0}\frac{k^{2}}{k_{S}^{2}}+a_{2}\mu^{2}\frac{k^{2}}{k_{S}^{2}}\right)\,,$ (13)
with $k_{S}=0.45~{}\,h\text{Mpc}^{-1}$. The addition of these parameters has minimal impact on the inference of cosmological parameters in the base $\Lambda$CDM model. Nonetheless, the EC pipeline incorporates both $\tilde{c}$ and $a_{0}$ contributions because it affects the error bars of cosmological parameters. The importance of the $k^{4}\mu^{4}P_{11}$ counterterm for consistency has been discussed in Ivanov et al. (2020a); Chudaykin et al. (2020, 2021b); Taule and Garny (2023); Ivanov et al. (2024b).
@@
$\begin{split}&\frac{c_{0}}{[\text{Mpc}/h]^{2}}\sim\mathcal{N}(0,10^{2})\,,\frac{c_{2}}{[\text{Mpc}/h]^{2}}\sim\mathcal{N}(0,60^{2})\,,\\ &a_{2}^{z_{1}}\sim\mathcal{N}(0,0.63^{2})\,,a_{2}^{z_{3}}\sim\mathcal{N}(0,0.44^{2})\\ &P_{\rm shot}^{z_{1}}\sim\mathcal{N}(0,0.11^{2})\,,P_{{\rm shot}}^{z_{3}}\sim\mathcal{N}(0,0.08^{2})\,,\\ &b_{1}\sim[0,4]_{\rm flat}\,,b_{2}\sim\mathcal{N}(0,5^{2})\,,b_{\mathcal{G}_{2}}\sim\mathcal{N}(0,0.4^{2})\,,\\ &b_{\Gamma_{3}}\sim\mathcal{N}\left(\frac{23}{42}(b_{1}-1),1^{2}\right)\,,\\ \end{split}$ (14)
where we approximated the boundaries in eq. (4) by Gaussian priors. Note that the stochastic contributions $P_{{\rm shot}}$ and $a_{2}$ are expressed in terms of the inverse mean galaxy density of the BOSS survey within the EC model.  [^7] The WC team uses a significantly narrower prior on $P_{\rm shot}$ with a standard deviation of $400\,(h^{-1}\text{Mpc})^{3}$ D’Amico et al. (2019). Surprisingly, this does not introduce a significant bias in the inference of cosmological parameters in the power-spectrum analyses of the $\Lambda$CDM model. The bias could be significant in extensions of this setup. In total, the WC model has 8 EFT parameters per sky chunk when fitting two multipoles.
Let us now specify the scale cuts in the WC model. The original analyses D’Amico et al. (2019); Colas et al. (2020) exploit a slightly different separation into LOWZ ($z_{\rm eff}=0.32$) and CMASS ($z_{\rm eff}=0.57$) redshift slices, and specify $k_{\rm max}^{\rm LOWZ}=0.2\,h\text{Mpc}^{-1}$ and $k_{\rm max}^{\rm CMASS}=0.23\,h\text{Mpc}^{-1}$, respectively. Since our $z_{3}$ slice has effectively a slightly higher redshift and a less data volume than CMASS, we re-scale the original $k_{\rm max}^{\rm CMASS}=0.23\,h\text{Mpc}^{-1}$ to $k_{\rm max}^{\rm z_{3}}=0.25\,h\text{Mpc}^{-1}$ using Eq. (40) of D’Amico et al. (2019). We keep $k_{\rm max}^{\rm z_{1}}=0.2\,h\text{Mpc}^{-1}$ following Simon et al. (2023).
@@
$P_{0}^{\text{stoch}.}\Big{|}_{\rm EC}=a_{2}\frac{k^{2}}{k_{\rm NL}^{2}}\frac{1}{3}\frac{1}{\bar{n}_{\rm EC}}\$ (15)
@@
$\begin{split}&\frac{c_{0}}{[\text{Mpc}/h]^{2}}\sim\mathcal{N}(0,30^{2})\,,\frac{c_{2}}{[\text{Mpc}/h]^{2}}\sim\mathcal{N}(30,30^{2})\,,\\ &\frac{\tilde{c}}{[\text{Mpc}/h]^{4}}\sim\mathcal{N}(0,500^{2})\,,P_{\rm shot}\sim\mathcal{N}(0,1^{2})\,,\\ &a_{0}\sim\mathcal{N}(0,1^{2})\,,a_{2}\sim\mathcal{N}(0,1^{2})\,,\\ &b_{1}\sim[0,4]_{\rm flat}\,,b_{2}\sim\mathcal{N}(0,1^{2})\,,b_{\mathcal{G}_{2}}\sim\mathcal{N}(0,1^{2})\,,\\ &b_{\Gamma_{3}}\sim\mathcal{N}\left(\frac{23}{42}(b_{1}-1),1^{2}\right)\,,\\ \end{split}$ (16)
@@
$\begin{split}&\frac{c_{4}}{[\text{Mpc}/h]^{2}}\sim\mathcal{N}(0,30^{2})\,,\frac{c_{1}}{[\text{Mpc}/h]^{2}}\sim\mathcal{N}(0,5^{2})\,,\\ &A_{\rm shot}\sim\mathcal{N}(0,1^{2})\,,B_{\rm shot}\sim\mathcal{N}(0,1^{2})\,.\\ \end{split}$ (17)
@@
Following the EC pipeline Philcox and Ivanov (2022), we change the scale cuts for the galaxy power spectrum to be $k_{\rm max}=0.2\,h\text{Mpc}^{-1}$ for the $z_{1}/z_{3}$ redshift bins. As we will see later, using a lower value of $k_{\rm max}^{z_{1}}$ effectively reduces the systematic bias in parameter recovery. We also adopt $k_{\rm max}=0.08\,h\text{Mpc}^{-1}$ for the bispectrum measurements.
@@ 4.1 PT Challenge analysis
Figure 1: Posterior distributions inferred from the single snapshot of the PT challenge simulation at $z_{\rm eff}=0.61$ in the WC and EC3 models using the analytical covariance matching the BOSS survey volume 5.8 ($h^{-1}$pc)3, and the covariance corresponding to the total simulation volume $566\,(h^{-1}\text{Gpc})^{3}$.
@@
$5.84\!\times\!V=$ 5.8 $V=5.8$ $V=566$ $5.84\!\times\!V=$ 5.8 $V=5.8$ $V=566$
$\Delta\omega_{cdm}/\omega_{cdm}$ $-0.038_{-0.033}^{+0.030}$ $-0.031_{-0.034}^{+0.032}$ $-0.015_{-0.0052}^{+0.0058}$ $-0.028_{-0.035}^{+0.031}$ $-0.015_{-0.037}^{+0.034}$ $-0.011_{-0.0050}^{+0.0045}$
$\Delta h/h$ $-0.004_{-0.012}^{+0.012}$ $-0.003_{-0.013}^{+0.012}$ $-0.0014_{-0.0014}^{+0.0015}$ $-0.002_{-0.013}^{+0.012}$ $-0.001_{-0.013}^{+0.012}$ $-0.0032_{-0.0014}^{+0.0014}$
$\Delta\sigma_{8}/\sigma_{8}$ $0.029_{-0.042}^{+0.040}$ $0.051_{-0.044}^{+0.041}$ $0.053_{-0.0051}^{+0.0049}$ $0.027_{-0.043}^{+0.040}$ $0.058_{-0.044}^{+0.041}$ $0.062_{-0.0046}^{+0.0045}$
$\Delta b_{1}/b_{1}$- $-0.077_{-0.046}^{+0.041}$ $-0.066_{-0.0062}^{+0.0065}$- $-0.069_{-0.046}^{+0.042}$ $-0.078_{-0.0049}^{+0.0050}$
$\Delta b_{2}$- $-0.77_{-0.61}^{+0.31}$ $-0.80_{-0.25}^{+0.21}$- $-0.38_{-0.62}^{+0.37}$ $-0.31_{-0.18}^{+0.17}$
$\Delta b_{\mathcal{G}_{2}}$- $0.34_{-0.29}^{+0.27}$ $-0.07_{-0.17}^{+0.15}$- $0.49_{-0.28}^{+0.27}$ $0.92_{-0.14}^{+0.14}$
@@
$5.84\!\times\!V=$ 5.8 $V=5.8$ $V=566$ $V=566$ $5.84\!\times\!V=$ 5.8 $V=5.8$ $V=566$
$\Delta\omega_{cdm}/\omega_{cdm}$ $-0.001_{-0.042}^{+0.039}$ $-0.006_{-0.048}^{+0.044}$ $-0.0085_{-0.0070}^{+0.0064}$ $-0.0066_{-0.0078}^{+0.0076}$ $0.002_{-0.034}^{+0.031}$ $0.010_{-0.037}^{+0.035}$ $0.0048_{-0.0046}^{+0.0043}$
$\Delta h/h$ $0.001_{-0.013}^{+0.013}$ $0.0_{-0.014}^{+0.013}$ $0.0002_{-0.0017}^{+0.0017}$ $-0.0034_{-0.0019}^{+0.0019}$ $0.001_{-0.013}^{+0.012}$ $0.001_{-0.013}^{+0.012}$ $0.0010_{-0.0013}^{+0.0013}$
$\Delta\sigma_{8}/\sigma_{8}$ $-0.018_{-0.046}^{+0.042}$ $0.037_{-0.047}^{+0.044}$ $0.051_{-0.0053}^{+0.0051}$ $0.026_{-0.0060}^{+0.0057}$ $-0.053_{-0.035}^{+0.031}$ $-0.013_{-0.042}^{+0.039}$ $0.014_{-0.0048}^{+0.0047}$
$\Delta b_{1}/b_{1}$- $-0.044_{-0.051}^{+0.049}$ $-0.057_{-0.0089}^{+0.0086}$ $-0.038_{-0.0087}^{+0.0091}$- $0.027_{-0.048}^{+0.046}$ $-0.0075_{-0.0059}^{+0.0057}$
$\Delta b_{2}$- $-0.33_{-0.80}^{+0.66}$ $-0.81_{-0.31}^{+0.29}$ $-0.62_{-0.43}^{+0.36}$- $0.11_{-0.41}^{+0.35}$ $-0.30_{-0.066}^{+0.064}$
$\Delta b_{\mathcal{G}_{2}}$- $0.25_{-0.37}^{+0.37}$ $-0.29_{-0.26}^{+0.26}$ $0.43_{-0.32}^{+0.32}$- $0.11_{-0.18}^{+0.17}$ $0.016_{-0.021}^{+0.020}$
Table 1: Marginalized 1d constraints on the cosmological and bias parameters inferred from the single- and multi-chunk analyses of the PT Challenge simulation data for different models. The cosmological parameters are shown as $(p-p_{\rm fid.})/p_{\rm fid.}$, where $p_{\rm fid.}$ denotes the fiducial value used in the simulation. For single-chunk analyses we also report constraints on the bias parameters. All volumes are expressed in units of $\,[h^{-1}\text{Gpc}]^{3}$.
@@
The magnitude of projection effects increases with the number of nuisance parameters, leading the single-chunk analysis to somewhat underestimate the parameter shifts due to marginalization effects. To reliable assess the prior volume effects, we perform a multi-chunk analysis of four simulation samples – two at the $z_{1}$ redshift bin ($z_{\rm eff}=0.38$) and two at the $z_{3}$ redshift bin ($z_{\rm eff}=0.61$). For each data chunk, we introduce an independent set of EFT parameters, allowing them to vary freely in the analysis. The mock catalogues at the same redshift slice are designed to be identical emulating the BOSS samples in the Northern and Southern galactic caps (NGC and SGC). We adjust the volume for each chunk to match the actual BOSS volumes of the NGCz1, SGCz1, NGCz3, and SGCz3 caps, while keeping the PT Challenge mean galaxy density $\bar{n}_{\rm PTC}$ to allow for meaningful comparisons with previous results. We call this analysis $5.84\times V=$ 5.8.

### v2

@@ Abstract
Parameter estimation from galaxy survey data from the full-shape method depends on scale cuts and priors on EFT parameters. The effects of priors, including the so-called “prior volume” phenomenon have been originally studied in Ivanov et al. (2019) and subsequent works. In this note, we repeat and extend these tests and also apply them to other priors used in the literature. We point out that in addition to the “prior volume” effect there is a more dangerous effect that is largely overlooked: a systematic bias on cosmological parameters due to overoptimistic scale cuts. Unlike the “prior volume” effect, this is a genuine systematic bias due to two-loop corrections that does not vanish with better priors or with larger data volumes. Our study is based on the high fidelity BOSS-like PT Challenge simulation data which offer many advantages over analyses based on synthetic data generated with fitting pipelines. We show that some analysis choices associated with the PyBird code, especially the scale cuts, significantly bias parameter recovery, overestimating $\sigma_{8}$ by over $5\%$ (equivalent to $1\sigma$). The bias on measured EFT parameters is even more significant. In contrast, the analysis choices associated with the CLASS-PT code lead to much smaller ($\lesssim 1\%$) shifts in cosmological parameters based on their best-fit values.
@@ 1 Introduction
The presence of the theory systematic error in EFT-FS analyses shows why parameter projection effects should not be considered in isolation. For instance, one may choose overoptimistic scale cuts or priors in their analysis, which would result in a very small prior volume effect, but both 1D posteriors and the best-fit values will be biased away from the true values. As we show here, certain analysis choices associated with the PyBird code D’Amico et al. (2021) lead to a non-negligible systematic bias in parameter recovery.
@@
The use of high-fidelity BOSS-like PT Challenge simulation data offers many advantages over analyses based on synthetic data generated within the fitting pipeline, done e.g. in Simon et al. (2023). First, it enables the estimation of the theory systematic bias arising from the imperfect modeling of the target statistics. Second, it allows one to estimate the quality of the recovery of true cosmological parameters. Third, the large-volume data allow one to easily estimate the true best-fit model by rescaling the covariance. This method is thus not sensitive to noise in the covariance matrices Hartlap et al. (2007); Sellentin and Heavens (2016); Howlett and Percival (2017); Wadekar et al. (2020); Philcox et al. (2021) and does not suffer from numerical uncertainties associated with the likelihood minimization.
@@ 2 Summary and Main Conclusions
We compare analysis choices associated with the PyBird and CLASS-PT EFT-FS likelihoods. Following Nishimichi et al. (2020) we dub those two analysis configurations as “West coast” (WC) and “East coast” (EC) models. We also consider two extensions of these models, EC1/2 designed to highlight certain analysis aspects. The EC model that matches the actual BOSS analysis Philcox and Ivanov (2022) is referred to as EC3. Note that the WC analysis here is the one was used in the actual BOSS analysis D’Amico et al. (2019) [^6] It is different from the WC analysis implemented in the PT Challenge Nishimichi et al. (2020), which used a more conservative $k_{\mathrm{max}}$ value.
@@
The recovered best-fit value of $\sigma_{8}$ is shifted up by $\approx 5\%$ as a result of the two-loop corrections. This bias remains in the combinations with other datasets like Planck CMB.
@@
Finally, our single-volume results can be treated as an approximation for ongoing stage-IV surveys such as DESI and Euclid, for which the volumes of individual data chunks will be comparable to the entire volume of BOSS DR12, c.f. Chudaykin and Ivanov (2019); Aghamousa et al. (2016) [^7] Notably, the DESI constraints are primarily informed by LRG galaxies, whose properties are very similar to those of the CMASS and LOWZ samples, so it is natural to extrapolate our results to DESI. In contrast, the ELG galaxies, one of the main targets of the Euclid mission, have a somewhat different morphology, so an accurate assessment of biases for ELGs requires a dedicated analysis. As we show here, in this situation the prior volume effects should not significantly complicate the interpretation of the results from these surveys.
@@ 3 Comparison of different analyses choices and priors
[…]
@@
where D’Amico et al. (2019) assumed $k_{M}=0.7\,h\text{Mpc}^{-1}$ and $P_{11}$ stands for the linear matter power spectrum at redshift $z$. In what follows we will suppress the explicit redshift dependence of $P_{11}$. Matching them, we find the following relation between $c_{0}$ and $c_{ct}$ parameters,
@@
where $k_{\rm NL}=0.45\,h\text{Mpc}^{-1}$, $\bar{n}_{\rm WC}$ and $\bar{n}_{\rm EC}$ refer the mean galaxy densities in the WC and EC models. Notice that WC assume $\bar{n}_{\rm WC}^{-1}=2500\,(h^{-1}\text{Mpc})^{3}$ and $2200\,(h^{-1}\text{Mpc})^{3}$ for the LOWZ and CMASS redshift bins, whereas in the EC pipeline $\bar{n}_{\rm EC}^{-1}=3500\,(h^{-1}\text{Mpc})^{3}$ and $5000\,(h^{-1}\text{Mpc})^{3}$ for $z_{1}$ and $z_{3}$ slices. Note that the actual measurements from data give $\bar{n}^{-1}=4800$ and 35003500 $(h^{-1}\text{Mpc})^{3}$ for CMASS and LOWZ, respectively Chen et al. (2024). The values of $\bar{n}$ for the $z_{1}/z_{3}$ data split in the EC pipeline correspond to the mean geometric number densities, i.e. the total number of galaxies divided by the effective geometric volume. These numbers are in good agreement with $\bar{n}(z_{\rm eff})$ measured directly from data using the appropriate weights and selection functions.
@@
Adopting $c_{\epsilon,\,\text{quad}}\sim\mathcal{N}(0,2^{2})$ from D’Amico et al. (2019), we observe that the WC prior on $a_{2}$ is tighter than that in the EC model which is $a_{2}\sim\mathcal{N}(0,1^{2})$. The inconsistency of this choice with the physics of the HOD-based galaxies was discussed in Ivanov et al. (2024b). As we will see later, this tight prior reduces the parameter shifts caused by marginalization effects in a multi-chunk full-shape EFT analysis. Importantly, the WC setup also excludes a scale-dependent stochastic counterterm in the monopole, $c_{\epsilon,\,\text{mono}}=0$. Although the real-space scale-dependent noise term can be set to zero based on the simulation results Schmittfull et al. (2019, 2020); Ivanov et al. (2024a, b), the contribution from the $k^{2}\mu^{2}$ counterterm is actually non-negligible, which raises concerns about the choice $c_{\epsilon,\,\text{mono}}=0$. Indeed, from the mathematical point of view, it is inconsistent to neglect a $k^{2}$ stochastic counterterm in the monopole if the $k^{2}$ stochastic contribution is present in the quadrupole, because physically the have the same origin in the $(k,\mu)$ space.
@@
$P^{\rm stoch}(k,\mu)=\frac{1}{\bar{n}}\left(1+P_{\rm shot}+a_{0}\frac{k^{2}}{k_{S}^{2}}+a_{2}\mu^{2}\frac{k^{2}}{k_{S}^{2}}\right)\,$, (13)
[…]
The importance of the $k^{4}\mu^{4}P_{11}$ counterterm for consistency has been discussed in Ivanov et al. (2020a); Chudaykin et al. (2020, 2021b); Taule and Garny (2023); Ivanov et al. (2024b). First, it is used to address the strong finger-of-the-God effect, which breaks the naive EFT counting and thus necessitates the inclusion of the next-to-leading order contribution. This term effectively captures the two-loop contribution induced by higher-derivative terms in the non-linear RSD mapping. As a result, its inclusion extends the validity of perturbation theory and reduces sensitivity to high-$k$ modes affected by the two-loop contributions. Second, dropping the $k^{4}\mu^{4}P_{11}$ term biases the inference of the EFT parameters, in particular $c_{\epsilon,~\text{quad}}~(a_{2})$, and $c_{r,1}~(c_{2})$ with which it is partly degenerate. While the $k^{4}\mu^{4}P_{11}$ and $k^{2}\mu^{2}$ terms are degenerate at the level of the monopole and quadrupole spectra, this degeneracy is broken by including the hexadecapole which allows one to measure both these terms independently. Thus, the $k^{4}\mu^{4}P_{11}$ term is important for the unbiased inference of both the leading-order counterterm coefficients and the $k^{2}\mu^{2}$ stochastic contribution.
@@
[…]
Let us now specify the scale cuts in the WC model. The original analyses D’Amico et al. (2019); Colas et al. (2020) exploit a slightly different separation into LOWZ ($z_{\rm eff}=0.32$) and CMASS ($z_{\rm eff}=0.57$) redshift slices, and specify $k_{\rm max}^{\rm LOWZ}=0.2\,h\text{Mpc}^{-1}$ and $k_{\rm max}^{\rm CMASS}=0.23\,h\text{Mpc}^{-1}$, respectively. Since our $z_{3}$ slice has effectively a slightly higher redshift and a less data volume than CMASS, we re-scale the original $k_{\rm max}^{\rm CMASS}=0.23\,h\text{Mpc}^{-1}$ to $k_{\rm max}^{\rm z_{3}}=0.25\,h\text{Mpc}^{-1}$ using the rescaling procedure suggested in Eq. (40) by D’Amico et al. (2019) for $z=0.61$. This rescaling has been used in many analyses. For instance, it has been employed in Simon et al. (2023). For that reason, we focus on $k_{\rm max}^{\rm z_{3}}=0.23\,h\text{Mpc}^{-1}$ in the main text. The analyses using a more conservative scale cut $k_{\rm max}^{\rm z_{3}}=0.23\,h\text{Mpc}^{-1}$ are presented in App. A. There we show that practically there is no difference between $k_{\rm max}^{\rm z_{3}}=0.23\,h\text{Mpc}^{-1}$ and $k_{\rm max}^{\rm z_{3}}=0.25\,h\text{Mpc}^{-1}$. We keep $k_{\rm max}^{\rm z_{1}}=0.2\,h\text{Mpc}^{-1}$ following Simon et al. (2023).
@@
$P_{0}^{\text{stoch}.}\Big|_{\rm EC}=a_{2}\frac{k^{2}}{k_{\rm NL}^{2}}\frac{1}{3}\frac{1}{\bar{n}_{\rm EC}}\\$ (15)
@@
[…]
@@
[…]
@@
Following the EC pipeline Philcox and Ivanov (2022), we change the scale cuts for the galaxy power spectrum to be $k_{\rm max}=0.2\,h\text{Mpc}^{-1}$ for the $z_{1}/z_{3}$ redshift bins. As we will see later, using a lower value of $k_{\rm max}^{z_{3}}$ effectively reduces the systematic bias in parameter recovery. We also adopt $k_{\rm max}=0.08\,h\text{Mpc}^{-1}$ for the bispectrum measurements.
@@ 4.1 PT Challenge analysis
Figure 1: Posterior distributions inferred from the single snapshot of the PT challenge simulation at $z_{\rm eff}=0.61$ in the WC and EC3 models using the analytical covariance matching the BOSS survey volume 5.8 ($h^{-1}$Gpc)3, and the covariance corresponding to the total simulation volume $566\,(h^{-1}\text{Gpc})^{3}$.
@@
$4\!\times\!V=5.8$ $V=5.8$ $V=566$ $4\!\times\!V=5.8$ $V=5.8$ $V=566$
[…]
$\Delta h/h$ $004_{-0.012}^{+0.012} 0.0$$03_{-0.013}^{+0.012} 0.0$$014_{-0.0014}^{+0.0015} 0.0$$02_{-0.013}^{+0.012} 0.0$$01_{-0.013}^{+0.012} 0.0$$0032_{-0.0014}^{+0.0014}$
$\Delta\sigma_{8}/\sigma_{8}$ $029_{-0.042}^{+0.040} 0.0$$51_{-0.044}^{+0.041} 0.0$$53_{-0.0051}^{+0.0049} 0.0$$27_{-0.043}^{+0.040} 0.0$$58_{-0.044}^{+0.041} 0.0$$062_{-0.0046}^{+0.0045}$
$\Delta b_{1}/b_{1}$ - $077_{-0.046}^{+0.041} 0.0$$66_{-0.0062}^{+0.0065} - 0.0$$69_{-0.046}^{+0.042} 0.0$$078_{-0.0049}^{+0.0050}$
$\Delta b_{2}$ - $0.77_{-0.61}^{+0.31}$ $0.80_{-0.25}^{+0.21} -$ $0.38_{-0.62}^{+0.37}$ $0.31_{-0.18}^{+0.17}$
$\Delta b_{\mathcal{G}_{2}}$ - $34_{-0.29}^{+0.27} 0.0$$07_{-0.17}^{+0.15} -$ $0.49_{-0.28}^{+0.27}$ $0.92_{-0.14}^{+0.14}$
@@
$4\!\times\!V=5.8$ $V=5.8$ $V=566$ $V=566$ $4\!\times\!V=5.8$ $V=5.8$ $V=566$
$\Delta\omega_{cdm}/\omega_{cdm}$ $001_{-0.042}^{+0.039} 0.0$$06_{-0.048}^{+0.044} 0.0$$085_{-0.0070}^{+0.0064} 0.0$$066_{-0.0078}^{+0.0076} 0.0$$02_{-0.034}^{+0.031} 0.0$$10_{-0.037}^{+0.035} 0.0$$0048_{-0.0046}^{+0.0043}$
$\Delta h/h$ $01_{-0.013}^{+0.013} 0.00.0$$_{-0.014}^{+0.013}$ $0002_{-0.0017}^{+0.0017} 0.0$$034_{-0.0019}^{+0.0019} 0.0$$01_{-0.013}^{+0.012} 0.0$$01_{-0.013}^{+0.012} 0.0$$0010_{-0.0013}^{+0.0013}$
$\Delta\sigma_{8}/\sigma_{8}$ $018_{-0.046}^{+0.042} 0.0$$37_{-0.047}^{+0.044} 0.0$$51_{-0.0053}^{+0.0051} 0.0$$26_{-0.0060}^{+0.0057} 0.0$$53_{-0.035}^{+0.031} 0.0$$13_{-0.042}^{+0.039} 0.0$$014_{-0.0048}^{+0.0047}$
$\Delta b_{1}/b_{1}$ - $044_{-0.051}^{+0.049} 0.0$$57_{-0.0089}^{+0.0086} 0.0$$38_{-0.0087}^{+0.0091} - 0.0$$27_{-0.048}^{+0.046} 0.0$$0075_{-0.0059}^{+0.0057}$
$\Delta b_{2}$ - $0.33_{-0.80}^{+0.66}$ $0.81_{-0.31}^{+0.29}$ $0.62_{-0.43}^{+0.36} -$ $0.11_{-0.41}^{+0.35}$ $0.30_{-0.066}^{+0.064}$
$\Delta b_{\mathcal{G}_{2}}$ - $0.25_{-0.37}^{+0.37}$ $0.29_{-0.26}^{+0.26}$ $0.43_{-0.32}^{+0.32}$ - $11_{-0.18}^{+0.17} 0.0$$016_{-0.021}^{+0.020}$
Table 1: Marginalized 1d constraints on the cosmological and bias parameters inferred from the single- and multi-chunk analyses of the PT Challenge simulation data for different models. The cosmological parameters are shown as $(p-p_{\rm fid.})/p_{\rm fid.}$, where $p_{\rm fid.}$ denotes the fiducial value used in the simulation. For single-chunk analyses we also report constraints on the bias parameters. All volumes are expressed in units of $h^{-1}\text{Gpc}]^{3}$.
@@
The magnitude of projection effects increases with the number of nuisance parameters, leading the single-chunk analysis to somewhat underestimate the parameter shifts due to marginalization effects. To reliable assess the prior volume effects, we perform a multi-chunk analysis of four simulation samples – two at the $z_{1}$ redshift bin ($z_{\rm eff}=0.38$) and two at the $z_{3}$ redshift bin ($z_{\rm eff}=0.61$). For each data chunk, we introduce an independent set of EFT parameters, allowing them to vary freely in the analysis. The mock catalogues at the same redshift slice are designed to be identical emulating the BOSS samples in the Northern and Southern galactic caps (NGC and SGC). We adjust the volume for each chunk to match the actual BOSS volumes of the NGCz1, SGCz1, NGCz3, and SGCz3 caps, while keeping the PT Challenge mean galaxy density $\bar{n}_{\rm PTC}$ to allow for meaningful comparisons with previous results. We call this analysis $4\times V=5.8$.